Could you please help me solve this system of equations? When I attempted to solve it I came out with (0,1,-5). This does not seem correct.
5x+3y+z=0
x-3y+2z=9
14x-2y+3z=11

Answer by oscargut(2103)

5x+3y+z=0
x-3y+2z=9
14x-2y+3z=11
5x+3y+z=0
9x+9y=-9
x+11y=-11
x+y=-1
x+11y=-11
x=-1-y
-1-y+11y=-11
10y=-10
y=-1
x=0
z=3
solution: (0,-1,3)